OpenStudy (maheshmeghwal9):

Prove that if \[x+\frac{1}{x}=2 \cos \alpha,\] then \[x^n+\frac{1}{x^n}= 2\cos n \alpha.\]. [This problem is based on DeMoivre's theorem]

OpenStudy (maheshmeghwal9):

I have done like this: - Let \[x= e^{i \theta}\]&\[\frac{1}{x}=e^{-i \theta}\]This approach is also given in my book. But now how to do further ? This is my actual problem.

OpenStudy (maheshmeghwal9):

@ujjwal @zepp @zzr0ck3r Plz help:)

OpenStudy (ujjwal):

\[x+\frac{1}{x}=2\cos \alpha\]only when x=1 and \(\alpha\)=0 This relation is not satisfied by any other values.

OpenStudy (ujjwal):

So, the second part goes accordingly!

OpenStudy (anonymous):

\[r\left(e^{i\theta} + e^{- i \theta}\right) = 2\cos \alpha\]

OpenStudy (ujjwal):

Contd. \(x^n\)=1 and n\(\alpha\)=0 for any value of zero Remember we already have x=1 and \(\alpha\)=0

OpenStudy (anonymous):

\[x^n = e^{ni\theta}\]

OpenStudy (maheshmeghwal9):

I gt that but isn't there logical way @ujjwal ?

OpenStudy (ujjwal):

*n In my last reply zero=n..

OpenStudy (anonymous):

\[x^n + \frac1{x^n} = e^{ni\theta}+e^{n-i\theta} = \cos n \theta + i\sin n\theta +\cos n \theta - i\sin n\theta = 2\cos n \theta \]

OpenStudy (ujjwal):

I know that is informal and there must be a very formal way to derive that relation.. And @Ishaan94 is giving it to you.

OpenStudy (maheshmeghwal9):

oh ok i see :) thanx to all for the help :)